Rosa purchased a sofa that was on sale for 20% off. She spent $1,850. How much was the sofa before sale?
Answer by ikleyn(52803)

.

    (1-0.2)x = 1850

       0.8x  = 1850

          x  = 1850/0.8 = 2312.50 dollars.     ANSWER

Solved.